# Simplifying Debiased Inference via Automatic Differentiation and Probabilistic Programming

### v1

@@ 2 $\frac{\gamma_{P}}{\sigma_{P}^{4}}$- $z\mapsto-\frac{[y-\mu_{P}(x)]^{2}-\gamma_{P}}{\sigma_{P}^{2}}$
In Appendix E, we apply Algorithm 2 to two other parameters: the expected density $\psi(P)=E_{P}[p(Z)]$ and expected conditional covariance $\psi(P)=E_{P}[\mathrm{cov}_{P}(A\,,Y\,|\,X)]$.
@@ 3.1 Estimator construction
Algorithm 3 presents a method to obtain this initial estimate and the evaluation of the efficient influence operator at a user-specified $f_{k}$. The algorithm employs a forward and backward pass. In the forward pass, it recursively estimates the $P$-dependent evaluations of the operators $\theta_{j}$ that appear in Algorithm 1, ultimately yielding the initial estimate $\widehat{h}_{k}$ of $\psi(P)$. In the backward pass, it recursively estimates $P$-dependent adjoint evaluations and backpropagates them as in Algorithm 2, yielding an estimate $\widehat{f}_{0}$ of the efficient influence operator evaluation $\dot{\psi}_{P}^{*}(f_{k})$. Black-box machine learning tools may be used to estimate all of the needed nuisances. When $\psi$ is real-valued and $f_{k}=1$, $\widehat{f}_{0}$ is an estimate of the EIF. To fix ideas, we focus on this real-valued setting in the remainder of the main text and refer the reader to Appendix C.3 for a discussion of more general settings where $\psi$ takes values in a generic separable Hilbert space, such as an $L^{2}$ space.
@@
and 95% confidence interval $\widehat{\psi}\pm 1.96\,\widehat{\sigma}/n^{1/2}$, where $\widehat{\sigma}^{2}$ is the empirical variance of $\widehat{f}_{0}$. In fact, $\widehat{\psi}$ is exactly the traditional one-step estimator $\psi(\widehat{P})+\frac{1}{n}\sum_{i=1}^{n}\dot{\psi}_{\widehat{P}}^{*}(1)(Z_{i})$ under the strong compatibility condition that there exists a $\widehat{P}\in\mathcal{M}$ such that $\widehat{h}_{j}=\theta_{j}(\widehat{P},\widehat{h}_{\mathrm{pa}(j)})$ and $\widehat{\vartheta}_{j}=\dot{\theta}_{j,\widehat{P},\widehat{h}_{\mathrm{pa}(j)}}^{*}(\widehat{f}_{j})$ for all $j\in[k]$. To ease implementation, Algorithm 3 makes a weaker compatibility requirement that allows each $\widehat{\vartheta}_{j,0}$, $\widehat{h}_{j}$, and $\widehat{h}_{\mathrm{pa}(j)}$ to be compatible with a $\widehat{P}_{j}$ that may depend on $j$. In the next subsection, we show that this allows each primitive to be paired with a nuisance estimation routine that can be reused across different parameters $\psi$.
@@
Our Python package implements a cross-fitted counterpart of $\widehat{\psi}$, which we present and study in Appendix C.1. Like $\widehat{\psi}$, this estimator is efficient when a remainder and drift term are negligible (see Theorem S1). The cross-fitted estimator may be preferred over $\widehat{\psi}$ because its drift term is $o_{p}(n^{-1/2})$ even without an empirical process condition (Schick, 1986). This is in contrast to $\widehat{\psi}$, for which such a condition is often used to ensure negligibility of that term (Lemma 19.24 of van der Vaart, 2000).
@@
where $\vartheta_{j,0}$ is the first entry of the augmentation term $\dot{\theta}_{j,P,h_{\mathrm{pa}(j)}}^{*}(f_{j})$ from Algorithm 2. If both terms on the right are $o_{p}(n^{-1/2})$, then $\mathcal{R}_{n}$ is as well. Moreover, since the above holds for any $\widehat{P}_{j}$ compatible with the nuisances $\widehat{h}_{j}$, $\widehat{h}_{\mathrm{pa}(j)}$, and $\widehat{\vartheta}_{j,0}$, it suffices that there exist such $\widehat{P}_{j}$, $j\in[k]$, that make each of the two terms on the right $o_{p}(n^{-1/2})$. By Cauchy-Schwarz, the magnitude of the latter term upper bounds by $\sum_{j=1}^{k}\|d\widehat{P}_{j}/dP-1\|_{L^{2}(P)}\|\widehat{\vartheta}_{j,0}-\vartheta_{j,0}\|_{L^{2}(P)}$, which is $o_{p}(n^{-1/2})$ under an $n^{-1/4}$-rate condition akin to those arising in the study of traditional one-step estimators. In Appendix C.2, we show that the first term on the right is the remainder in a von Mises expansion, which makes it so it too will be $o_{p}(n^{-1/2})$ under appropriate conditions. Similar arguments have been made previously to motivate one-step estimators when $\psi$ is a generic pathwise differentiable parameter (Robins et al., 2009; Luedtke and Chung, 2023).
For a particular parameter $\psi$, the remainder $\mathcal{R}_{n}$ can also be analyzed explicitly. In Appendix F, we provide this analysis for the nonparametric $R^{2}$ parameter.
@@ 3.2 Software implementation
The dimple package also adopts computation graphs, using a static graph structure (Baydin et al., 2018). When defining the parameter, each primitive call takes dimple objects as inputs and returns another dimple object. At this stage, a dimple object comprises a primitive and a set of pointers to its inputs, serving the roles of $\theta_{j}$ and $\mathrm{pa}(j)$ in Algorithm 1. The dimple objects can be used to express $\psi$ as in Algorithm 1, and also provide the form of the computation graph. Estimation — as executed, for instance, via ‘estimate(R2)’ in Example LABEL:ex:r2 — calls Algorithm 3 with this computation graph to populate each dimple object with initial estimates $\widehat{h}_{j}$ in the forward pass, with adjoint estimates $\widehat{f}_{j}$ added in the backward pass.
@@
Forward and backward routines for the primitives in Table 1 are given in Appendix G. Those routines allow needed regression, density functions, and density ratios to be estimated using flexible statistical learning tools, such as gradient-boosted trees (Ke et al., 2017). Riesz losses (Chernozhukov et al., 2022c) are used when estimating the adjoint of bounded affine maps whose ambient Hilbert spaces depend on $P$.
4 Discussion
Our results suggest researchers should refocus their efforts when constructing an efficient estimator of a new parameter. Traditionally, this process involves labor-intensive calculations to derive the EIF. Our approach avoids this task for parameters expressible as compositions of known primitives. Yet, we recognize that this method will not apply to all new parameters, particularly those that cannot be represented in this form. In such cases, we propose a shift in focus: rather than immediately studying the differentiability of the parameter itself, researchers should aim to express it as a composition involving both known and novel primitives. By establishing the differentiability of each new primitive, they can deduce the differentiability of the parameter itself. This modular approach can both simplify the immediate analysis and the study of other parameters in future research.
As noted in Carone et al. (2018), any automated approach to statistical inference inevitably has potential for misuse. Per Theorems 1 and 2, the validity of our method depends on differentiability and nuisance estimation conditions. While an expert can interpret and verify these conditions, non-specialists may employ a software implementation without doing so. This may lead to an overconfidence in the inferential conclusions reached, especially in high-dimensional problems or other settings where these conditions are unlikely to be met. To mitigate this risk, future software could automatically provide a list of conditions sufficient for ensuring the differentiability of each primitive and the rates of convergence required on nuisance estimators for efficiency. Though Appendix D already gives sufficient conditions for the differentiability of the primitives in Table 1, work is still needed to figure out how to have software return this information in an accessible manner.
There are also other interesting avenues for future work. First, though our algorithms apply to arbitrary models, the particular nuisance estimation routines given in Appendix G are designed for nonparametric settings. Future work should identify semiparametric model restrictions that allow for modular nuisance estimation. Second, it would be interesting to consider whether it is possible to automatically modify a representation of a parameter to weaken the conditions required of nuisance estimators. For example, suppose $\psi(P)=E_{P}[Z]$ is expressed as $\langle z\mapsto z,dP/d\lambda\rangle_{L^{2}(\lambda)}$ using inner product and density primitives. Algorithm 3 would estimate the nuisance $dP/d\lambda$ for this representation and the efficiency of the one-step estimator would rely on an $n^{-1/4}$ rate condition. Estimation of this nuisance, and the associated rate condition, would be avoided entirely if $\psi$ were instead expressed using a marginal mean primitive. Automatically finding alternative representations of a parameter that weaken nuisance estimation requirements is an important area for future work.
Acknowledgements
This work was supported by the National Science Foundation and National Institutes of Health under award numbers DMS-2210216 and DP2-LM013340. The author is grateful to Jonathan Kernes for kindly letting him use code from his Medium tutorial in the dimple package (Kernes, 2021). He is also grateful to Saksham Jain for taking the time to test the package on a sunny Seattle afternoon.
Appendices
[…]
1
Appendix A Simulation study
@@
A simulation study was conducted to evaluate the performance of the dimple package for the three examples from Section 1.3. For a given Monte Carlo replicate, a dataset dat was generated as an iid sample of $n\in\{250,1000,4000,16000\}$ draws from a distribution $P$, and then the dimple code from Section 1.3 was executed to perform estimation. A total of 1000 Monte Carlo replicates were performed for each scenario. Code for this study is available at http://github.com/alexluedtke12/dimple.
@@
For the expected density parameter, we followed Carone et al. (2018) in letting $P=\mathrm{Beta}(3,5)$. For the nonparametric R-squared parameter, we followed Williamson et al. (2021) in letting $P$ be the distribution of $(X,Y)$ with $X=(X_{1},X_{2})$ a pair of independent $\mathrm{Unif}[-1,1]$ random variables and $Y\,|\,X\sim N(25X_{1}^{2}/9,1)$. For the longitudinal G-formula example, we used the data-generating process from Simulation 1 of Luedtke et al. (2017).
@@
Table S1 reports the results. The coverage of 95% confidence intervals returned by dimple approximately attains the nominal value for all scenarios with $n\geq 1000$. Coverage tends to be below nominal when $n=250$, taking a minimal value of 79%. The widths of the confidence intervals are close to the asymptotically optimal width of $2\times 1.96\times\sigma/n^{1/2}$, with $\sigma$ the standard deviation of the EIF. For each expected density and nonparametric R-squared scenario, the variance of the estimators always approximately equals the efficient variance, $\sigma^{2}/n$. In the longitudinal G-formula scenario, this variance is much larger than the efficient variance at smaller sample sizes but the relative difference diminishes as $n$ grows. The inflated variance appears to be partially driven by outliers arising from inverse propensity weighting in the EIF. Bias is minimal in all scenarios. Together these results support our theoretical findings that dimple can produce asymptotically efficient estimators under reasonable conditions (Theorem 2).
@@
Appendix B Total pathwise differentiability: supplemental theoretical results and proofs
B.1   Preliminary lemmas
@@ B.3   Proof of Theorem 1
B.4   Relationship between total pathwise differentiability and other notions of differentiability
@@
We call $\dot{\zeta}_{u}$ the differential operator of $\zeta$ at $u$. As noted in Section 20.7 of van der Vaart (2000), Hadamard differentiability is equivalent to Gateaux differentiability uniformly over compacts. Moving now to pathwise differentiability (van der Vaart, 1991a; Luedtke and Chung, 2023), $\nu:\mathcal{M}\rightarrow\mathcal{W}$ is called pathwise differentiable at $P\in\mathcal{M}$ if there exists a continuous linear operator $\dot{\nu}_{P}:\dot{\mathcal{M}}_{P}\rightarrow\mathcal{W}$ such that, for all $s\in\check{\mathcal{M}}_{P}$ and $\{P_{\epsilon}:\epsilon\}\in\mathscr{P}(P,\mathcal{M},s)$,
@@
Appendix C Estimation: supplemental theoretical results and proofs
C.1   Theoretical guarantees

### v3

@@ 2 $\frac{\gamma_{P}}{\sigma_{P}^{4}}$- $z\mapsto-\frac{[y-\mu_{P}(x)]^{2}-\gamma_{P}}{\sigma_{P}^{2}}$
In Appendix D, we apply Algorithm 2 to two other parameters: the expected density $\psi(P)=E_{P}[p(Z)]$ and expected conditional covariance $\psi(P)=E_{P}[\mathrm{cov}_{P}(A\,,Y\,|\,X)]$.
@@ 3.1 Estimator construction
Algorithm 3 presents a method to obtain this initial estimate and the evaluation of the efficient influence operator at a user-specified $f_{k}$. The algorithm employs a forward and backward pass. In the forward pass, it recursively estimates the $P$-dependent evaluations of the operators $\theta_{j}$ that appear in Algorithm 1, ultimately yielding the initial estimate $\widehat{h}_{k}$ of $\psi(P)$. In the backward pass, it recursively estimates $P$-dependent adjoint evaluations and backpropagates them as in Algorithm 2, yielding an estimate $\widehat{f}_{0}$ of the efficient influence operator evaluation $\dot{\psi}_{P}^{*}(f_{k})$. Black-box machine learning tools may be used to estimate all of the needed nuisances. When $\psi$ is real-valued and $f_{k}=1$, $\widehat{f}_{0}$ is an estimate of the EIF. To fix ideas, we focus on this real-valued setting in the remainder of the main text and refer the reader to Appendix B.3 for a discussion of more general settings where $\psi$ takes values in a generic separable Hilbert space, such as an $L^{2}$ space—these more general settings cover, for example, the case where $\psi$ returns a square integrable dose-response function..
@@
and 95% confidence interval $\widehat{\psi}\pm 1.96\,\widehat{\sigma}/n^{1/2}$, where $\widehat{\sigma}^{2}$ is the empirical variance of $\widehat{f}_{0}$. In fact, $\widehat{\psi}$ is exactly the traditional one-step estimator $\psi(\widehat{P})+\frac{1}{n}\sum_{i=1}^{n}\dot{\psi}_{\widehat{P}}^{*}(1)(Z_{i})$ under the strong compatibility condition that there exists a $\widehat{P}\in\mathcal{M}$ such that $\widehat{h}_{j}=\theta_{j}(\widehat{P},\widehat{h}_{\mathrm{pa}(j)})$ and $\widehat{\vartheta}_{j}=\dot{\theta}_{j,\widehat{P},\widehat{h}_{\mathrm{pa}(j)}}^{*}(\widehat{f}_{j})$ for all $j\in[k]$. To ease implementation, Algorithm 3 makes a weaker compatibility requirement that allows each $\widehat{\vartheta}_{j,0}$, $\widehat{h}_{j}$, and $\widehat{h}_{\mathrm{pa}(j)}$ to be compatible with a $\widehat{P}_{j}$ that may depend on $j$. This dependence makes it so $\widehat{\psi}$ may not be a traditional one-step estimator; the benefit of permitting this is that it allows each primitive to be paired with a nuisance estimation routine that can be implemented once and reused across different parameters $\psi$—see the next subsection for details.
@@
Our Python package implements a cross-fitted counterpart of $\widehat{\psi}$, which we present and study in Appendix B.1. Like $\widehat{\psi}$, this estimator is efficient when a remainder and drift term are negligible (see Theorem S1). The cross-fitted estimator may be preferred over $\widehat{\psi}$ because its drift term is $o_{p}(n^{-1/2})$ even without an empirical process condition (Schick, 1986). This is in contrast to $\widehat{\psi}$, for which such a condition is often used to ensure negligibility of that term (Lemma 19.24 of van der Vaart, 2000).
@@
where $\vartheta_{j,0}$ is the first entry of the augmentation term $\dot{\theta}_{j,P,h_{\mathrm{pa}(j)}}^{*}(f_{j})$ from Algorithm 2. If both terms on the right are $o_{p}(n^{-1/2})$, then $\mathcal{R}_{n}$ is as well. Moreover, since the above holds for any $\widehat{P}_{j}$ compatible with the nuisances $\widehat{h}_{j}$, $\widehat{h}_{\mathrm{pa}(j)}$, and $\widehat{\vartheta}_{j,0}$, it suffices that there exist such $\widehat{P}_{j}$, $j\in[k]$, that make each of the two terms on the right $o_{p}(n^{-1/2})$. By Cauchy-Schwarz, the magnitude of the latter term upper bounds by $\sum_{j=1}^{k}\|d\widehat{P}_{j}/dP-1\|_{L^{2}(P)}\|\widehat{\vartheta}_{j,0}-\vartheta_{j,0}\|_{L^{2}(P)}$, which is $o_{p}(n^{-1/2})$ under an $n^{-1/4}$-rate condition akin to those arising in the study of traditional one-step estimators. In Appendix B.2, we show that the first term on the right is the remainder in a von Mises expansion, which makes it so it too will be $o_{p}(n^{-1/2})$ under appropriate conditions. Similar arguments have been made previously to motivate one-step estimators when $\psi$ is a generic pathwise differentiable parameter (Robins et al., 2009; Luedtke and Chung, 2024).
For a particular parameter $\psi$, the remainder $\mathcal{R}_{n}$ can also be analyzed explicitly. In Appendix E, we provide this analysis for the nonparametric $R^{2}$ parameter, and we also provide an analogue of Table 2 that helps to visualize Algorithm 3.
@@ 3.2 Software implementation
Our proof-of-concept package also adopts computation graphs, using a static graph structure (Baydin et al., 2018). When defining the parameter, each primitive call takes dimple objects as inputs and returns another dimple object. At this stage, a dimple object comprises a primitive and a set of pointers to its inputs, serving the roles of $\theta_{j}$ and $\mathrm{pa}(j)$ in Algorithm 1. The dimple objects can be used to express $\psi$ as in Algorithm 1, and also provide the form of the computation graph. Estimation — as executed, for instance, via ‘estimate(R2)’ in Example LABEL:ex:r2 — calls Algorithm 3 with this computation graph to populate each dimple object with initial estimates $\widehat{h}_{j}$ in the forward pass, with adjoint estimates $\widehat{f}_{j}$ added in the backward pass.
@@
Forward and backward routines for the primitives in Table 1 are given in Appendix F. Those routines allow needed regression, density functions, and density ratios to be estimated using flexible statistical learning tools, such as gradient-boosted trees (Ke et al., 2017). Riesz losses (Chernozhukov et al., 2022c) are used when estimating the adjoint of bounded affine maps whose ambient Hilbert spaces depend on $P$.
4 Simulation study
@@
A simulation study was conducted to verify whether a handful of lines of pydimple code truly suffice to construct performant one-step estimators for the examples in Section 1.3. Such a finding would be important given that existing implementations of one-step estimators in these problems can consist of hundreds of lines of code (e.g., Luedtke et al., 2017; Williamson et al., 2021). For a given Monte Carlo replicate of our study, a dataset dat was generated as an iid sample of $n\in\{250,1000,4000,16000\}$ draws from a distribution $P$, and then the pydimple code from Section 1.3 was executed to perform estimation. A total of 1000 Monte Carlo replicates were performed for each scenario. Code for this study is available at http://github.com/alexluedtke12/pydimple.
@@
For the expected density parameter, we followed Carone et al. (2018) in letting $P=\mathrm{Beta}(3,5)$. For the nonparametric $R^{2}$ parameter, we followed Williamson et al. (2021) in letting $P$ be the distribution of $(X,Y)$ with $X=(X_{1},X_{2})$ a pair of independent $\mathrm{Unif}[-1,1]$ random variables and $Y\,|\,X\sim N(25X_{1}^{2}/9,1)$. For the longitudinal G-formula example, we used the data-generating process from Simulation 1 of Luedtke et al. (2017).
@@
Table 3 reports the results. The coverage of 95% confidence intervals returned by pydimple approximately attains the nominal value for all scenarios with $n\geq 1000$. Coverage tends to be below nominal when $n=250$, taking a minimal value of 79%. The widths of the confidence intervals are close to the asymptotically optimal width of $2\times 1.96\times\sigma/n^{1/2}$, with $\sigma$ the standard deviation of the EIF. For each expected density and nonparametric $R^{2}$ scenario, the variance of the estimators always approximately equals the efficient variance, $\sigma^{2}/n$. In the longitudinal G-formula scenario, this variance is much larger than the efficient variance at smaller sample sizes but the relative difference diminishes as $n$ grows. The inflated variance appears to be partially driven by outliers arising from inverse propensity weighting in the EIF. Bias is minimal in all scenarios. Together these results support our theoretical findings that dimple can produce asymptotically efficient estimators under reasonable conditions (Theorem 2).
@@
5 Discussion
Our results suggest researchers should refocus their efforts when constructing an efficient estimator of a new parameter. Traditionally, this process involves labor-intensive calculations to derive the EIF. Our approach avoids this task for parameters expressible as compositions of known primitives. Yet, we recognize that this method will not apply to all new parameters, particularly those that cannot be represented in this form. In such cases, we propose a shift in focus: rather than immediately studying the differentiability of the parameter itself, researchers should aim to express it as a composition involving both known and novel primitives. By establishing the differentiability of each new primitive, they can deduce the differentiability of the parameter itself. This modular approach can both simplify the immediate analysis and the study of other parameters in future research. An example of a primitive that would be useful for ongoing research—but is missing from Table 1—is one that returns a solution to an infinite-dimensional inverse problem (Bennett et al., 2023; Cui et al., 2024; Smucler et al., 2025).
As noted in Carone et al. (2018), any algorithmic approach to statistical inference inevitably has potential for misuse. Per Theorems 1 and 2, the validity of our method depends on differentiability and nuisance estimation conditions. While an expert can interpret and verify these conditions, non-specialists may employ a software implementation without doing so. This may lead to an overconfidence in the inferential conclusions reached, especially in high-dimensional problems or other settings where these conditions are unlikely to be met. To mitigate this risk, future software could provide a list of conditions sufficient for ensuring the differentiability of each primitive and the rates of convergence required on nuisance estimators for efficiency. Though Appendix C already gives sufficient conditions for the differentiability of the primitives in Table 1, work is still needed to figure out how to have software return this information in an accessible manner.
The composition of primitives used to express the parameter $\psi$ will generally be nonunique. The same holds in traditional automatic differentiation settings, where, for example, the function $f(x)=2x^{2}+1$ can be expressed as $g\circ h$ with $g(x)=2x+1$ and $h(x)=x^{2}$, $g(x)=\cosh(x)$ and $h(x)=2\sinh^{-1}(x)$, or $g(x)=2x^{1/2}+1$ and $h(x)=x^{4}$. In traditional settings, choosing different compositions can impact the conditions under which the chain rule can be applied—for example, $g(x)=2x^{1/2}+1$ is not differentiable at 00, even though $f(x)=2x^{2}+1$ is. As indicated by the differentiability conditions in Theorem 1, the applicability of the chain rule also depends on the choice of representation for $\psi$ in our statistical setting. Beyond that, this choice can also impact the nuisances that need to be estimated. For example, suppose $\psi(P)=E_{P}[Z]^{2}$ is expressed as $\langle z\mapsto z,dP/d\lambda\rangle_{L^{2}(\lambda)}^{2}$ using inner product and density primitives. Algorithm 3 would estimate the nuisance $dP/d\lambda$ for this representation and the efficiency of the one-step estimator would rely on an $n^{-1/4}$ rate condition. Estimation of this nuisance, and the associated rate condition, would be avoided if $\psi$ were instead expressed using a marginal mean primitive and square function. Algorithmically finding alternative representations of a parameter that weaken nuisance estimation requirements is an important area for future work.
There are also other interesting avenues for future work. First, though our algorithms apply to arbitrary models, the particular nuisance estimation routines given in Appendix F are designed for nonparametric settings. Future work should seek to identify semiparametric model restrictions that allow for modular nuisance estimation. Second, while we focused on algorithmically constructing one-step estimators, it would be worth considering whether this can be done for other estimation strategies, such as targeted minimum loss-based estimation (van der Laan and Rubin, 2006; van der Laan and Gruber, 2016). Third, it would be interesting to investigate whether automatic differentiation can be used to obtain higher-order influence functions (Robins et al., 2008).
Acknowledgements
This work was supported by the National Science Foundation and National Institutes of Health under award numbers DMS-2210216 and DP2-LM013340. The author is grateful to Jonathan Kernes for kindly letting him use code from his Medium tutorial in pydimple (Kernes, 2021). He is also grateful to Saksham Jain for taking the time to test the package on a sunny Seattle afternoon.
Appendices
[…]
1
Appendix A Total pathwise differentiability: supplemental theoretical results and proofs
A.1   Preliminary lemmas
@@ A.3   Proof of Theorem 1
A.4   Relationship between total pathwise differentiability and other notions of differentiability
@@
We call $\dot{\zeta}_{u}$ the differential operator of $\zeta$ at $u$. As noted in Section 20.7 of van der Vaart (2000), Hadamard differentiability is equivalent to Gateaux differentiability uniformly over compacts. Moving now to pathwise differentiability (van der Vaart, 1991a; Luedtke and Chung, 2024), $\nu:\mathcal{M}\rightarrow\mathcal{W}$ is called pathwise differentiable at $P\in\mathcal{M}$ if there exists a continuous linear operator $\dot{\nu}_{P}:\dot{\mathcal{M}}_{P}\rightarrow\mathcal{W}$ such that, for all $s\in\check{\mathcal{M}}_{P}$ and $\{P_{\epsilon}:\epsilon\}\in\mathscr{P}(P,\mathcal{M},s)$,
@@
Appendix B Estimation: supplemental theoretical results and proofs
B.1   Theoretical guarantees